OpenStudy (anonymous):

x 2x - 4 ___ - _______ = 5 3 2 x 2x - 4 ___ - _______ = 5 3 2 @Mathematics

OpenStudy (anonymous):

The cost of Hisako's school supplies was $49.80. She had a coupon for 30% off the entire purchase. What was the final price?

OpenStudy (anonymous):

Get everything over a common denomiator then multiply through by it. 2x-6x+12 = 30 Re-arrange and solve for x

OpenStudy (anonymous):

multiply thru by 6 2x - 3(2x-4) = 30 -4x = 18 x = -4.5